### v1

# Hao Ma1, Zhiyuan Peng2, Xu Li3, Yukai Li1, Mingjie Shao1, Qiuqiang Kong4, and Ju Liu1

Abstract
Language-queried target sound extraction (TSE) aims to extract specific sounds from mixtures based on language queries. Traditional fully-supervised training schemes require extensively annotated parallel audio-text data, which are labor-intensive. We introduce a language-free training scheme, requiring only unlabelled audio clips for TSE model training by utilizing the multi-modal representation alignment nature of the contrastive language-audio pre-trained model (CLAP). In a vanilla language-free training stage, target audio is encoded using the pre-trained CLAP audio encoder to form a condition embedding for the TSE model, while during inference, user language queries are encoded by CLAP text encoder. This straightforward approach faces challenges due to the modality gap between training and inference queries and information leakage from direct exposure to target audio during training. To address this, we propose a retrieval-augmented strategy. Specifically, we create an embedding cache using audio captions generated by a large language model (LLM). During training, target audio embeddings retrieve text embeddings from this cache to use as condition embeddings, ensuring consistent modalities between training and inference and eliminating information leakage. Extensive experiment results show that our retrieval-augmented approach achieves consistent and notable performance improvements over existing state-of-the-art with better generalizability.
Index Terms:
target sound extraction, uni-modal training for multi-modal tasks, contrastive language-audio pre-training, retrieval-augmented strategy
I Introduction
Figure 1: Overview of our proposed retrieval-augmented language-free training scheme for language-queried target sound extraction.
Humans possess a remarkable ability to focus on specific sounds in noisy environments, a phenomenon known as the cocktail party effect. In signal processing, sound separation [1, 2] has been extensively researched to address this challenge. Early work focused on domain-specific signals like speech [3, 4] or music [5, 6], but recent advances in deep learning have led to universal sound separation (USS) [7], with the ambition to generalize the sound separation model to all kinds of sounds. A key development in USS is framing it as language-queried target sound extraction (TSE) [8, 9, 10, 11], where natural language descriptions are used to extract specific sounds from audio mixtures.
Despite significant progress in language-queried TSE, practical challenges persist. Current TSE systems [8, 9, 11] heavily depend on large-scale text annotations of sound events for training. This fully-supervised approach requires extensive parallel audio-text datasets to enhance model accuracy and generalizability. However, the high labor cost of annotations limits the availability of parallel data, making it insufficient for scalable universal sound extraction. This scarcity, though under-explored in TSE, is a known challenge in other cross-modal tasks like text-to-image [12], text-to-audio [13], and audio captioning [14]. These studies [12, 13, 14] address the issue by leveraging modality-aligned embedding spaces from contrastive cross-modal pre-trained models, eliminating the need for extensive parallel data.
Building on this approach, the contrastive language-audio pre-trained model (CLAP) [15] can be used to eliminate the need for text annotations in TSE. Typically, TSE models are trained using a target audio-text pair mixed with non-target audio, learning to extract the target sound in the audio mixture based on the text query. To bypass text annotations, CLAP can be cascaded with the TSE model. During training, the target audio is encoded into an audio embedding by the CLAP audio encoder, which the TSE model uses to extract the target sound. During testing, a text query is encoded by the CLAP text encoder into a text embedding for sound extraction. However, this vanilla approach assumes perfect alignment between text and audio embeddings, which is unrealistic. Two major challenges arise: the persistent modality gap between text and audio, worsened by training-testing inconsistencies, and the risk of overfitting due to the exposure of rich auditory details in the target audio embedding during training.
To address these challenges, we propose a retrieval-augmented, language-free training paradigm for language-queried TSE models as detailed in Fig. 1. Before training, we use a large language model to generate diverse audio captions, which are encoded into text embeddings by the pre-trained CLAP text encoder and stored in an embedding cache. During training, we retrieve the most similar text embeddings based on the target audio embedding and use them as the condition embedding, guiding the TSE model in sound extraction. This retrieval approach mitigates the modality mismatch and information leakage issues. Additionally, we apply noise injection [12, 16, 14] to the conditional embedding as an augmentation strategy, enhancing model generalizability. Extensive experiments show that our retrieval-augmented approach significantly improves performance and generalizability compared to existing state-of-the-art methods. By relaxing the need for parallel data, our method scales easily for large-scale training and outperforms previous supervised training schemes across multiple benchmarks, achieving a 1-2 dB improvement in signal-to-distortion ratio (SDR).
II Language-queried Target Sound Extraction
In this section, we propose to address the challenging task of language-queried target sound extraction. We first present the traditional fully-supervised training scheme when parallel audio-text data is available. Then, we introduce the proposed language-free training scheme with no need for parallel audio-text data.
II-A Fully-Supervised Training with Parallel Data
In fully-supervised training, we have access to audio-text pairs ${\bm{d}}=\{{\bm{x}},{\bm{y}}\}$, where ${\bm{x}}$ is an audio clip and ${\bm{y}}$ is its corresponding annotation text. In the mix-and-separate training pipeline, the audio mixture is constructed by sampling several audio clips, treating one of them as the target source and the others as noise, and then summing them as:
$\tilde{{\bm{x}}}={\bm{x}}+{\bm{v}},$ (1)
where $\tilde{{\bm{x}}}\in{\mathbb{R}}^{N}$ denotes the sound mixture of length $N$, ${\bm{x}}\in{\mathbb{R}}^{N}$ denotes the target sound source, and ${\bm{v}}\in{\mathbb{R}}^{N}$ denotes other interfering components. A query-conditioned target sound extraction model $\mathcal{F}(\cdot)$ parameterized by $\theta$ learns a map from the sound mixture to the target source conditioned on the target source information as:
$\hat{{\bm{x}}}=\mathcal{F}(\tilde{{\bm{x}}},{\bm{c}};\theta),$ (2)
where $\hat{{\bm{x}}}\in{\mathbb{R}}^{N}$ represents the target sound source prediction and ${\bm{c}}\in{\mathbb{R}}^{D}$ denotes the $D$-dimensional condition embedding, which is acquired by encoding the annotation text of the target sound by the CLAP text encoder both in the training and inference stages, which is:
${\bm{c}}=\text{CLAP}_{\text{text}}({\bm{y}}).$ (3)
II-B Proposed Approach
In language-free training without parallel data, we only have access to audio clips without text annotations. The critical challenge is how to construct the condition embedding ${\bm{c}}$ without access to text annotations? Our methods take advantage of the modality-aligned embedding space of contrastive language-audio pre-trained models. In the following part of this section, we first present a vanilla language-free training scheme and the widely adopted Gaussian noise injection as an augmentation strategy. We then propose a retrieval-augmented strategy to further mitigate specific challenges posed by existing methods to achieve improved performance.
II-B1 Vanilla Language-Free Training
In the vanilla language-free training, the condition embedding ${\bm{c}}$ is constructed by directly encoding the target audio ${\bm{x}}$ with the pre-trained CLAP audio encoder as:
${\bm{c}}=\text{CLAP}_{\text{audio}}({\bm{x}}).$ (4)
During inference, we use the pre-trained CLAP text encoder to encode the user language query to control the TSE model. This vanilla language-free training scheme works thanks to the modality-aligned embedding space generated by pre-trained CLAP, but it also suffers from two performance degradation points: one is the modality gap arose from the imperfect alignment of audio and text embeddings [17, 18]; the other is the information leakage by directly exposing the target audio during training that makes the model prone to overfitting.
II-B2 Gaussian Noise Injection
Gaussian noise injection has been proven to be effective in bridging the modality gap in many other cross-modal tasks [12, 16, 14]. This method models the gap between the audio and text embeddings extracted by pre-trained CLAP encoders on parallel audio and text as Gaussian noise with mean 00 and variance $\epsilon$. Based on the above modeling, the Gaussian noise is used to perturb the condition embeddings extracted from the target audio during language-free training of a TSE model as:
${\bm{c}}=\text{CLAP}_{\text{audio}}({\bm{x}})+{\bm{n}},$ (5)
where ${\bm{n}}\in{\mathbb{R}}^{D}$ is sampled from a Gaussian distribution $N(\mathbf{0},\epsilon{\bm{I}})$. The noise variance $\epsilon$ is an important hyperparameter in training, which will be explored in Section III-C.
II-B3 Retrieval-Augmented Strategy
Although the modality gap is effectively addressed by Gaussian noise injection, the issue of information leakage remains a point of concern. We further propose a retrieval-augmented strategy to address both the modality gap and the information leakage problem simultaneously. As shown in Fig. 1, before the model training, we first generate massive audio captions by prompting a large language model following [19]. These audio captions are then encoded by the pre-trained CLAP text encoder to create an embedding cache ${\bm{E}}\in{\mathbb{R}}^{M\times D}$, where $M$ is the number of audio captions generated by LLM.
Algorithm 1 retrieval-augmented language-free training scheme.
1:Audio dataset with $K$ audio samples $\mathcal{X}=\{{\bm{x}}_{1},{\bm{x}}_{2},...,{\bm{x}}_{K}\}$, embedding cache ${\bm{E}}\in\mathbb{R}^{M\times D}$, randomly initialized TSE model $\mathcal{F}(\tilde{{\bm{x}}},{\bm{c}};\theta)$
2:Optimized model parameters $\theta^{*}$
3:Initialize model parameters $\theta$
4:repeat
5:    Sample a mini-batch of audio: $\mathcal{X}_{i}\leftarrow\mathrm{sample}(\mathcal{X})\in\mathbb{R}^{B\times N}$;
6:    Create audio mixture: $\mathcal{\tilde{X}}_{i}\leftarrow\mathcal{X}_{i}+\mathrm{shuffle}(\mathcal{X}_{i})\in\mathbb{R}^{B\times N}$;
7:    Extract target audio embedding: ${\bm{q}}\leftarrow\text{CLAP}_{\text{audio}}(\mathcal{X}_{i})\in\mathbb{R}^{B\times D}$;
8:    // Retrieve from the embedding cache
9:    Compute similarity: $\mathrm{sim}\leftarrow{\bm{q}}\cdot{\bm{E}}^{T}\in\mathbb{R}^{B\times M}$;
10:    Find indices of maximum similarity:
11:    $\mathrm{indices}\leftarrow\mathrm{argmax}(\mathrm{sim},\text{axis}=1)\in\mathbb{R}^{B}$;
12:    Retrieve text embeddings: ${\bm{c}}\leftarrow{\bm{E}}[\mathrm{indices},:]\in\mathbb{R}^{B\times D}$;
13:    Noise injection: ${\bm{c}}\leftarrow{\bm{c}}+{\bm{n}}_{1}\sim N(\mathbf{0},\epsilon_{1}{\bm{I}})\in\mathbb{R}^{B\times D}$;
14:    Compute the gradient: $\nabla_{\theta}\mathcal{L}(\mathcal{X}_{i},\mathcal{F}(\mathcal{\tilde{X}}_{i},{\bm{c}};\theta))$;
15:    Update model parameters $\theta$ based on the computed gradient;
16:until model convergence
17:return optimized $\theta^{*}\leftarrow\theta$
TABLE I: Laguage-queried TSE performance evaluation. “bal.” denotes AudioSet balanced_train and “unbal.” denotes AudioSet unbalanced_train. The term “Vanilla” denotes target audio embeddings w/o any augmentation as condition embeddings for training and “NI” is noise injection for short. Except for LASS and AudioSep, all the methods listed in this table are built upon CLAPSep with different training strategies. †: reproduction in [11], ‡: “others” in [9] is comprised of AudioCaps, Clotho v2, WavCaps, VGGSound.
Method Parallel Dataset AudioCaps Clotho v2 AudioSet MUSIC21 ESC50 Avg
Data SDRi SI-SDRi SDRi SI-SDRi SDRi SI-SDRi SDRi SI-SDRi SDRi SI-SDRi SDRi SI-SDRi
LASS [8] Y AudioCaps 7.37 6.06 5.93 3.01 5.00 2.02 2.13 -3.64 8.32 6.00 5.75 2.69
AudioSep [9] Y unbal.+others 8.12 7.03 7.26 5.64 8.44 7.26 9.22 7.85 10.23 9.21 8.65 7.40
CLAPSep [11] Y AudioCaps 9.69 8.78 8.63 6.80 8.52 6.88 5.86 1.90 10.88 9.23 8.72 6.72
weakly-supervised Y bal. 7.76 6.33 6.95 3.93 8.13 6.69 7.65 5.91 10.55 9.23 8.21 6.42
-w/ NI Y bal. 7.92 6.58 7.81 5.82 8.04 6.67 7.77 6.27 10.79 9.87 8.47 7.04
Vanilla N bal. 7.33 5.81 6.57 4.65 4.33 -0.13 5.31 2.15 8.20 6.15 6.35 3.73
-w/ NI N bal. 8.65 7.62 8.38 6.85 6.63 3.99 8.02 6.29 10.99 10.15 8.53 6.98
Retrieval N bal. 8.76 7.73 8.40 6.73 6.78 4.11 8.23 6.59 11.26 10.40 8.69 7.11
-w/ NI N bal. 8.70 7.67 8.60 7.14 7.17 4.87 8.28 6.78 11.36 10.63 8.82 7.42
Vanilla w/ NI N $\mathcal{DS}_{large}$ 9.60 8.71 9.06 7.59 7.11 4.20 9.85 8.37 12.20 11.40 9.56 8.05
Retrieval w/ NI N $\mathcal{DS}_{large}$ 9.75 8.92 9.43 8.12 8.09 5.75 10.24 9.11 12.55 11.89 10.00 8.76
In training, the target audio ${\bm{x}}$ is encoded by the CLAP audio encoder as a query to retrieve the most similar text embedding in the embedding cache ${\bm{E}}$. The retrieved text embedding is considered as a condition embedding ${\bm{c}}$ and further augmented by a Gaussian perturbation as:
${\bm{c}}=\mathrm{Retrieve}({\bm{E}},\text{query}=\text{CLAP}_{\text{audio}}({\bm{x}}))+{\bm{n}}_{1},$ (6)
where the retrieval function is detailed in Alg. 1 and ${\bm{n}}_{1}\sim N(\mathbf{0},\epsilon_{1}{\bm{I}})\in{\mathbb{R}}^{D}$. We explore the setting of $\epsilon_{1}$ in Section III-C.
By retrieving pre-encoded text embeddings, we address the modality gap because the queries are in the same text modality in both the training and inference phases. Meanwhile, by using retrieved condition embeddings, the TSE model does not directly see the target audio embeddings during its training, further avoiding information leakage. A more detailed procedure for retrieval-augmented language-free training of a language-queried target sound extraction model is given in Alg. 1.
III Experiments
III-A Datasets and Evaluation Metrics
We consider two data sources commonly used in the literature [9, 11, 10, 7] for model training: AudioSet [20] and FreeSound [21]. AudioSet is a large-scale audio collection from YouTube videos, while FreeSound is an online collaborative sound-sharing site. We collect a total of 1,912,173 audio clips from the unbalanced_train split of AudioSet and 262,300 audio clips from FreeSound. All audio clips are sampled at 32kHz.
Firstly, we perform quick preliminary experiments on the balanced_train split of AudioSet, which contains 20,550 audio clips, to evaluate the effectiveness of our proposed method and determine the hyperparameter settings. We then scale up the training data to include all collected audio clips from AudioSet and FreeSound to further validate the effectiveness of our method on large-scale non-parallel data. We refer to this large dataset as $\mathcal{DS}_{large}$.
We follow recipes in [9, 11] to prepare evaluation datasets. We perform evaluations on AudioCaps [22], Clotho v2 [23], the AudioSet test split, MUSIC21 [24] and ESC50 [25] to comprehensively validate the effectiveness of our proposed method. Note that in AudioCaps and Clotho v2, the annotations are human-written audio captions in natural language form, while in AudioSet, MUSIC21, and ESC50, the annotations are labels of sound events. In testing, we convert these labels to language-like audio captions by adding the prefix “The sound of ”.
We adopt the commonly used SDRi and SI-SDRi [26] as evaluation metrics, which are defined as:
$\mathrm{SDRi}(\hat{{\bm{x}}},\tilde{{\bm{x}}},{\bm{x}})$ $=\mathrm{SDR}(\hat{{\bm{x}}},{\bm{x}})-\mathrm{SDR}(\tilde{{\bm{x}}},{\bm{x}}),$ (7)
$\mathrm{SI\text{-}SDRi}(\hat{{\bm{x}}},\tilde{{\bm{x}}},{\bm{x}})$ $=\mathrm{SI\text{-}SDR}(\hat{{\bm{x}}},{\bm{x}})-\mathrm{SI\text{-}SDR}(\tilde{{\bm{x}}},{\bm{x}}),$ (8)
where
$\mathrm{SDR}(\hat{{\bm{x}}},{\bm{x}})$ $=10*\mathrm{log}_{10}\left(\frac{||{\bm{x}}||^{2}}{||{\bm{x}}-\hat{{\bm{x}}}||^{2}}\right),$ (9)
$\mathrm{SI\text{-}SDR}(\hat{{\bm{x}}},{\bm{x}})$ $=10*\mathrm{log}_{10}\left(\frac{||\frac{||\hat{{\bm{x}}}^{T}{\bm{x}}||^{2}}{||{\bm{x}}||^{2}}{\bm{x}}||^{2}}{||\frac{||\hat{{\bm{x}}}^{T}{\bm{x}}||^{2}}{||{\bm{x}}||^{2}}{\bm{x}}-\hat{{\bm{x}}}||^{2}}\right).$ (10)
III-B Implementation Details
We build the language-free training framework upon the state-of-the-art (SOTA) language-queried target sound extraction model, CLAPSep [11]. For experiments on the balanced_train split, we train the model on a single RTX 3090 GPU with a batch size of 32 for a total of 96,450 steps. The initial learning rate is set to 1e-4, which decays exponentially by a factor of 0.32 at steps 31,150 and 64,300. For experiments on $\mathcal{DS}_{large}$, we train the model on four RTX 3090 GPUs with a global batch size of 64 for a total of 509,400 steps. The initial learning rate is set to 2e-4, which decays exponentially by a factor of 0.32 at steps 169,800 and 339,600. We adopt the negative SI-SDR between the model estimation and the target audio as an end-to-end loss function as detailed in equation 10.
III-C Results and Analysis
In this section, we first perform preliminary experiments on AudioSet balanced_train to demonstrate the validity of the proposed method and to determine the hyperparameter settings. Then we scale up the training data to demonstrate the effectiveness of our proposed method in leveraging large-scale unlabelled audio data for training a language-queried TSE model with strong generalizability.
Effectiveness of retrieval-augmented strategy. We show the performance of various language-free training schemes in Table I. As can be seen from these results, the vanilla strategy, i.e., the direct use of target audio embedding without any augmentation as the condition embedding for training, performs the worst. This indicates that the modality gap and information leakage introduced by this strategy cause a significant drop in the performance of the TSE model. The modality gap problem can be solved to some extent by Gaussian noise injection. As can be seen from the results in the table (Vanilla w/ NI on bal.), Gaussian noise injection brings more than 2dB performance improvement on average for the vanilla strategy. Further, the retrieval-augmented strategy proposed in this paper combined with a certain degree of noise injection (Retrieval w/ NI on bal.) performs optimally over all kinds of language-free training schemes. It is worth noting that the better performance achieved by our proposed method does not rely on Gaussian noise injection, which is shown by the performance gap between “Retrieval” and “Retrieval w/ NI” is smaller than that between vanilla strategies. This indicates that the proposed method has effectively addressed the modal gap and information leakage problems by retrieving pre-encoded text embeddings.
Language-free versus weakly-supervised. We conduct weakly-supervised experiments by invoking the label-formed annotations on AudioSet and transforming the audio labels into captions by prefixing each with “The sound of ”. The detailed results on each of the evaluation benchmarks are shown in Table I. Comparing the weakly-supervised training scheme with our proposed language-free training scheme, the weakly-supervised method suppresses the language-free methods only on AudioSet but falls behind on other evaluation benchmarks. This is because the query language in the AudioSet test set is defined in the same way as in weakly-supervised training. This phenomenon shows that our proposed approach has better generalizability to various query forms, especially those in natural language form, as demonstrated by the results on AudioCaps and Clotho v2.
Ablation study of Gaussian perturbation strength. The choice of the Gaussian perturbation variance has a significant effect on the performance of language-free training. To determine the Gaussian perturbation strength, we vary the variation of injected Gaussian noise from 0 to 1e-1 and train a TSE model for each value. We run evaluations on all five testing datasets and present the results of averaged SDRi and SI-SDRi in Fig. 2.
Figure 2: Effect of Gaussian perturbation strength.
The green lines in Fig. 2 illustrate the effect of noise variance $\epsilon$ on the performance of the standard Gaussian noise injection strategy, which injects Gaussian noise directly into the target audio embedding to form the condition embedding in training. As shown in Fig. 2, the appropriate choice of noise variance has a crucial impact on the effectiveness of this approach. Specifically, when the noise variance is small, the modality gap and information leakage will cause a significant drop in performance. On the other hand, if the noise variance is too large, the useful information in audio embeddings is overwhelmed by too much noise. According to Fig. 2, we set the noise variance to 1e-2 for all subsequent experiments for the standard Gaussian noise injection strategy.
The red lines in Fig. 2 illustrate the effect of the noise variance $\epsilon_{1}$ on the performance of our proposed retrieval-augmented strategy. The experimental results show that for our proposed retrieval-augmented strategy, Gaussian noise injection of appropriate intensity also improves performance to some extent, but the effect is far less dramatic than for the standard noise injection strategy, further demonstrating that the better performance achieved by the proposed retrieval-augmented strategy does not depend on the noise perturbation.
Ablation study of LLM-produced captions. Our proposed retrieval-augmented language-free training scheme requires retrieving corresponding text embeddings from a pre-encoded text embedding cache to construct condition embeddings. Therefore, the quality of this embedding cache plays a crucial role in the performance of the proposed retrieval augmented strategy. We suggest using LLM-produced audio captions to construct this embedding cache. In this section, we design experiments to verify the impact of LLM-produced audio captions on the performance of our proposed method. Specifically, we first establish a baseline system where we take all the 527 audio labels from AudioSet and construct 527 audio captions by prefixing each with “The sound of ”. These captions are then encoded into text embeddings using the CLAP text encoder and stored in an embedding cache, thus forming our baseline system. Then we invoke WavCaps [19] to simulate LLM-produced audio captions. We encode all the 403,050 LLM-produced audio captions from WavCaps into text embeddings, which are additionally added to the embedding cache in the baseline system. We train a new TSE model for each of the two settings and perform evaluation experiments on all five datasets. The results are shown in Fig. 3. The experimental results show that the proposed method, combined with the LLM-produced audio captions, achieves significant performance improvements on most evaluation benchmarks.
Scale up the training data. In this section, we scale up the training data to include all collected audio clips from AudioSet and FreeSound to further validate the effectiveness of our method in leveraging large-scale non-parallel data. The corresponding results are presented in the last two rows of Table I. From these results, we can see that the proposed language-free training scheme performs better on all evaluation benchmarks after scaling up the training data. These results outperform the current SOTA supervised training schemes on most of the evaluated benchmarks, thus demonstrating the effectiveness of the proposed method in efficiently leveraging a large amount of audio data without annotation to train a language-queried target sound extraction model with strong generalizability.
Figure 3: Ablation study of LLM-produced captions.
IV Conclusion
In this work, we present a novel language-free training paradigm for language-queried target sound extraction, leveraging the modality-aligned embedding space generated by the pre-trained CLAP model. By harnessing unlabeled audio clips, our model deftly sidesteps the necessity for labor-intensive text-audio pairings. We address challenges like modality gap and information leakage through a retrieval-augmented strategy, employing an embedding cache constructed from LLM-produced audio captions. Our extensive evaluation on several benchmarks demonstrates that this language-free training approach achieves consistent and notable performance improvements over existing state-of-the-art with better generalizability. This method significantly reduces the reliance on annotated data, offering a more adaptable and scalable solution for TSE.